Write equation in slope-intercept form for the line that passes through (5, 0) and is perpendicular to the line described by y=-5/2x +6
Answer by lmeeks54(105)

The slope-intercept form of an equation is:
y = mx + b, where:
m = the slope of the line
b = the y intercept (where x = 0)
...
Given:
y = -5/2 x + 6
line passes through the point (5, 0). Note points are in the form (x, y)
...
problem: find a line parallel to line given above. Parallel means has to have the same slope.
...
When we have a point, like (5, 0), that means that point is a solution to the equation. We already know m (slope), x, and y, we need to solve for b to create our parallel line:
...
Given:
m = -5/2 (or -2.5)
x = 5
y = 0
Find b
...
0 = -5/2 (5) + b
0 = -12.5 + b
...
add 12.5 to both sides:
12.5 = b
...
Solution equation is:
y = -2.5 x + 12.5
...
cheers,
Lee